OpenStudy (anonymous):

2x over 5 + 3x over 2?

OpenStudy (anonymous):

Multiply both sides by 2 to get rid of the fraction there... 2 {3x / 2 = x + 15} 3x = 2x + 30 3x - 2x = 30 x = 30

OpenStudy (callisto):

\[\frac{2x}{5} + \frac{3x}{2}=\frac{2x(2)}{5(2)} + \frac{3x(5)}{2(5)} = \frac{4x+15x}{10} = \frac{19x}{10}\]

OpenStudy (callisto):

I thought it was addition of fraction... @krizzypoo12 Is your question 'solve the equation' or 'simplify the expression'?

OpenStudy (anonymous):

perform the addition or subtraction indicated..